Understanding Gas Laws:


Part B
A gas, behaving ideally, fills a fixed volume container at a pressure P1 and at a temperature T1. The temperature of the container is changed to T2. Using Avogadro’s, Charles’, or Boyle’s law, develop an expression that would solve for the new pressure P2.
A gas, behaving ideally, fills a fixed volume container at a pressure and at a temperature . The temperature of the container is changed to . Using Avogadro’s, Charles’, or Boyle’s law, develop an expression that would solve for the new pressure .
A: P1/T1 = P2/T2
B: P1V1 = P2V2
C: P1T1 = P2T2
D: V1/T1 = V2/T2
1
Expert's answer
2017-03-22T09:48:46-0400
Question:
Understanding Gas Laws:
Part B
A gas, behaving ideally, fills a fixed volume container at a pressure P1 and at a temperature T1. The temperature of the container is changed to T2. Using Avogadro’s, Charles’, or Boyle’s law, develop an expression that would solve for the new pressure P2.
A gas, behaving ideally, fills a fixed volume container at a pressure and at a temperature . The temperature of the container is changed to . Using Avogadro’s, Charles’, or Boyle’s law, develop an expression that would solve for the new pressure .
A: P1/T1 = P2/T2
B: P1V1 = P2V2
C: P1T1 = P2T2
D: V1/T1 = V2/T2

Solution:
I have to note that required relation can not be developed via three mentioned laws. It is independent expression commonly known as the Gay-Lussac’s law (but I have to recognize that in some sources the same expression is referred to as the Charle’s law) and it looks the following way:
P/T=k, where
P - is the pressure of the gas,
T - is the temperature of the gas,
k - is a constant;

From here if system changes it’s state from state 1 to state 2, than
P1/T1= P2/T2 (=k);

and the final expression for the new pressure is following:
P2 = P1T2 / T1.

Answer:
A: P1/T1 = P2/T2

The final expression is:
P2 = P1T2 / T1.